Martha can rake the leaves in her yard in 2 hours. Her yonger brother can do the job in 3 hours. How long will it take them to do the job if thy work together?
A)5/6hr
B)6hr
C)3hr
D)6/5hr
and why?

Answer by Nate(3500)

Martha's Rate: 1/2
Martha's Brother's Rate: 1/3
Rate * Time = Work
t/2 + t/3 = 1
6t/2 + 6t/3 = 6
3t + 2t = 6
5t = 6
t = 6/5
(D)